CF/(1+r) - INV vestment the questions 1. Based on the following information, answer the questions, 1) Facebook issued 10-yearPlease help with part 1, #1,2,3
Answer #1

1.
The bond price is less than 1000 because coupon rate is less than yield to maturity

If coupon rate is more than yield to maturity, bond price is more than par value
If coupon rate is less than yield to maturity, bond price is less than par value
If coupon rate is equal to yield to maturity, bond price is equal to par value

2.
N=10*2=20
PMT=-8%*1000/2=-40
FV=-1000
I/Y=9%/2=4.5%
CPT PV=934.960318

3.
N=10*2=20
PMT=8%*1000/2=40
PV=-990
FV=1000
CPT I/Y=4.07407%

Yield to maturity=4.07407%*2=8.14814%